# Geometry Induced Chiral Transport and Entanglement in $AdS_{2}$ Background

### v1

@@ 2.1 Fermions on Two-Dimensional Curved Spaces
We will work in two-dimensional curved space-time with metric $g_{\mu\nu}$, related to the flat Minkowski metric $\eta_{ab}$ by the vielbein $e_{\mu}^{a}$ by:
@@
where $e_{\mu}^{a}e_{b}^{\mu}=\delta^{a}_{b}$. We consider a massive Dirac fermion, whose action reads:
@@
where the adjoint spinor is defined as $\bar{\psi}=\psi^{\dagger}\gamma^{0}$, where $\gamma^{0}$ is the flat space gamma matrix, and $A=(A_{t},0)$ is a time-like vector field. $D_{\mu}$ is the fermionic covariant derivative  [^1] The reason for the $i$ multiplying the spin connection term is that in our notations $\gamma_{0}\gamma_{1}$ and $\omega$ are both real, hence this term is hermitian.:
$D_{\mu}\psi=\partial_{\mu}\psi+\frac{i}{4}\omega_{\mu}^{ab}\gamma_{a}\gamma_{b}\psi\ ,$ (4)
where the spin connection $\omega_{\mu}^{ab}$ is defined as:
$\omega_{\mu}^{ab}=e^{a\nu}\left(\partial_{\mu}e^{b}_{\nu}-\partial_{\nu}e^{b}_{\mu}\right)+e^{a\nu}e^{b\sigma}\Gamma_{\mu\nu}^{\rho}e_{\rho c}\eta^{bc}\ .$ (5)
@@ 2.2.1 Action and Hamiltonian
The Lagrangian density takes the form:
${\cal L}=\psi^{\dagger}\left[i\frac{1}{\sqrt{f(r})}\partial_{t}+i\sqrt{f(r)}\sigma_{x}\partial_{r}-\frac{r}{2L^{2}\sqrt{f(r)}}\sigma_{x}-m\sigma_{z}+\mu\right]\psi\ .$ (11)
@@
${\cal L}=\chi^{\dagger}\left[i\partial_{t}+if(r)\sigma_{x}\partial_{r}-\frac{r}{2L^{2}}\sigma_{x}-\sqrt{f(r)}m\sigma_{z}+\sqrt{f(r)}\mu\right]\chi\ ,$ (15)
@@
$H=\int_{r_{h}}^{\infty}dr\mathcal{H}=\int dr\chi^{\dagger}\left[-if(r)\sigma_{x}\partial_{r}+\frac{r}{2L^{2}}\sigma_{x}+\sqrt{f(r)}m\sigma_{z}-\sqrt{f(r)}\mu\right]\chi\ .$ (16)
@@
We will consider the weighted charge that multiplies the chemical potential $\mu$ in the Hamiltonian (16):
@@ 2.2.2 Qubit Hamiltonian
Upon the regularization, our Hamiltonian with mass and chemical potential is Ikeda:2025yqq
@@
The first term in the Hamiltonian is identical to the kinetic term in flat space. The subsequent term involves the spin connection, represented by $\frac{1}{2z}\psi^{\dagger}\sigma_{x}\psi=\frac{1}{2an}\bar{\psi}\gamma^{1}\psi$. The mass term and the chemical potential are $L$-dependent. If we take the limit $L\to\infty$ while maintaining $\frac{L}{an}=1$, the Hamiltonian reduces to the flat space.

### v2

@@ 2.1 Fermions on Two-Dimensional Curved Spaces
We will work in two-dimensional curved space-time with metric $g_{\mu\nu}$, related to the flat Minkowski metric $\eta_{ab}=\mathrm{diag}(-,+)$ by the vielbein $e_{\mu}^{a}$ by:
@@
where $e_{\mu}^{a}e_{b}^{\mu}=\delta^{a}_{b}$. The flat-frame gamma matrices satisfy the two-dimensional Clifford algebra:
$\{\gamma^{a},\gamma^{b}\}=-2\,\eta^{ab}\ ,$ (3)
and we will use the representation:
$\gamma^{0}=\sigma_{z},\gamma^{1}=i\sigma_{y},\gamma^{5}=\gamma^{0}\gamma^{1}=\sigma_{x}\ ,$ (4)
and $(\gamma^{0})^{\dagger}=\gamma^{0},(\gamma^{1})^{\dagger}=-\gamma^{1},(\gamma^{5})^{\dagger}=\gamma^{5}$.
We consider a massive Dirac fermion, whose action reads:
@@
where the adjoint spinor is defined as $\bar{\psi}=\psi^{\dagger}\gamma^{0}$, where $\gamma^{0}$ is the flat space gamma matrix, and $A=(A_{t},0)$ is a time-like vector field. $D_{\mu}$ is the fermionic covariant derivative:
$D_{\mu}\psi=\partial_{\mu}\psi-\frac{1}{4}\omega_{\mu}^{ab}\gamma_{a}\gamma_{b}\psi\ ,$ (6)
where the torsionless spin connection $\omega^{a}{}_{b}=\omega^{a}{}_{b\mu}dx^{\mu}$ satisfies the Cartan equation:
$de^{a}+\omega^{a}{}_{b}\wedge e^{b}=0\ ,$ (7)
where $e^{a}=e^{a}_{\mu}dx^{\mu}$.
@@ 2.2.1 Action and Hamiltonian
The Lagrangian density takes the form: [^1] The operator $\frac{i}{2}\{A(r),\partial_{r}\}\;=\;i\,A(r)\,\partial_{r}\;+\;\frac{i}{2}\,A^{\prime}(r)$ is hermitian. Here $A(r)=\sqrt{f(r)}\,\sigma_{x}$.
${\cal L}=\psi^{\dagger}\left[i\frac{1}{\sqrt{f(r})}\partial_{t}+i\sqrt{f(r)}\sigma_{x}\partial_{r}+i\frac{r}{2L^{2}\sqrt{f(r)}}\sigma_{x}-m\sigma_{z}+\mu\right]\psi\ .$ (13)
@@
${\cal L}=\chi^{\dagger}\left[i\partial_{t}+if(r)\sigma_{x}\partial_{r}+i\frac{r}{L^{2}}\sigma_{x}-\sqrt{f(r)}m\sigma_{z}+\sqrt{f(r)}\mu\right]\chi\ ,$ (17)
@@
$H=\int_{r_{h}}^{\infty}dr\mathcal{H}=\int dr\chi^{\dagger}\left[-if(r)\sigma_{x}\partial_{r}-i\frac{r}{L^{2}}\sigma_{x}+\sqrt{f(r)}m\sigma_{z}-\sqrt{f(r)}\mu\right]\chi\ .$ (18)
@@
We will consider the weighted charge that reflects the underlying curved spacetime geometry:
@@ 2.2.2 Qubit Hamiltonian
Upon the regularization, our Hamiltonian with mass is Ikeda:2025yqq
@@
The first term in the Hamiltonian is identical to the kinetic term in flat space. The subsequent term involves the spin connection, represented by $\frac{1}{2z}\psi^{\dagger}\sigma_{x}\psi=\frac{1}{2an}\bar{\psi}\gamma^{1}\psi$. The mass term is $L$-dependent. If we take the limit $L\to\infty$ while maintaining $\frac{L}{an}=1$, the Hamiltonian reduces to the flat space.